Using matrix method, solve the simultaneous equations

{x-3y=3

5x-9y=11


1
Expert's answer
2021-09-06T19:20:18-0400
"AX=B"

"A^{-1}AX=A^{-1}B"

"X=A^{-1}B"

"A=\begin{pmatrix}
 1 & -3 \\
 5 & -9
\end{pmatrix}, X=\begin{pmatrix}
 x \\
 y
\end{pmatrix},B=\begin{pmatrix}
 3 \\
 11
\end{pmatrix}"

"\begin{pmatrix}
 1 & -3 \\
 5 & -9
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}=\begin{pmatrix}
 3 \\
 11
\end{pmatrix}"

"\det A=\begin{vmatrix}
 1 & -3 \\
 5 & -9
\end{vmatrix}=1(-9)-(-3)(5)=6\not=0"

"A^{-1}=\dfrac{1}{6}\begin{pmatrix}
 -9 & 3 \\
 -5 & 1
\end{pmatrix}=\begin{pmatrix}
 -3/2 & 1/2 \\
 -5/6 & 1/6
\end{pmatrix}"

"\begin{pmatrix}
 -3/2 & 1/2 \\
 -5/6 & 1/6
\end{pmatrix}\begin{pmatrix}
 3\\
 11
\end{pmatrix}=\begin{pmatrix}
 -9/2+11/2 \\
 -15/6+11/6
\end{pmatrix}"

"=\begin{pmatrix}
 1 \\
 -2/3
\end{pmatrix}"

"x=1, y=-\dfrac{2}{3}"

"(1, -\dfrac{2}{3})"